OpenStudy (torii897):

Jose bought a wallet for $19.95 and y hats for $5.99 each. Which of the following expressions shows the total amount of money he spent? A: 19.95y + 5.99y B:19.95y + 5.99 C:19.95 + 5.99y D:19.95 + 5.99 + y

Mehek (mehek14):

which one do you think it is?

Mehek (mehek14):

@torii897 ?

OpenStudy (torii897):

D?

Mehek (mehek14):

no

Mehek (mehek14):

each hat (y) costs 5.99 so to find how much you pay for y number of hats, you would multiply 5.99y = 5.99 * y

Mehek (mehek14):

so which one would it be?

Mehek (mehek14):

ok just leave without answering

OpenStudy (torii897):

I don't understand this its hard

OpenStudy (torii897):

@Mehek14

Mehek (mehek14):

ok look at what I said \(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 each hat (y) costs 5.99 so to find how much you pay for y number of hats, you would multiply 5.99y = 5.99 * y \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (torii897):

I multiplied it but I got 35.8801

OpenStudy (torii897):

@Mehek14

Mehek (mehek14):

why would you multiply????

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 ok look at what I said \(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 each hat (y) costs 5.99 so to find how much you pay for y number of hats, you would multiply 5.99y = 5.99 * y \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\)

Mehek (mehek14):

so part of your problem has 5.99y

Mehek (mehek14):

and that person only buys 1 wallet for 19.95

Mehek (mehek14):

C:19.95 + 5.99y

OpenStudy (torii897):

okay thanks for helping me im not that smart I don't get math @Mehek14

Mehek (mehek14):

it's fine :) it'll take some time to learn but you will get it eventually :)